A current of 9.65 ampere is passed through the aqueous solution of NaCl using suitable electrodes for 1000s.The amount of NaOH formed during electrolysis?

Grade:12th pass

1 Answers

ishita
6 years ago
The liberation of each h+ ion can result in 1 naoh being formed .moles of electron formed = (9.65×1000)÷96500 = 0.1so moles of h+ liberated =0.1 moles of naoh = 0.1 mass of naoh formed = 40×0.1= 4.0 g